in a triangle the measure of the first angle is three times the measure of the second angle. the measure of the third angle is 20 degrees less than the second angle. what is the measure of each angle

Answer by checkley79(3341)

LET 3X BE THE FIRST ANGLE
LET X BE THE SECOND ANGLE.
LET X-20 BE THE THIRD ANGLE.
3X+X+(X-20)=180
5X=180+20
5X=200
X=200/5
X=40 DEGREES FOR THE SECOND ANGLE
3*40=120 DEGREES FOR THE FIRST ANGLE.
40-20=20 DEGREES FOR THE THIRD ANGLE.
PROOF:
120+40+20=180
180=180